# custom_references.bib

### v1

@@ Abstract
In fault-tolerant distributed computing, two fundamental tasks are collective coin flipping—where processors agree on a common random bit—and leader election—where they designate a leader among themselves. We study these problems in the full-information model, where processors communicate via a single broadcast channel, have access to private randomness, and face a computationally unbounded adversary that controls some of the processors. Despite decades of study, key gaps remain in our understanding of the trade-offs between round complexity, communication per player in each round, and adversarial resilience.
We make progress by proving new lower bounds for coin flipping protocols, which also imply lower bounds for leader election protocols. Specifically, we show that any $k$-round coin flipping protocol, where each of $\ell$ players sends $1$ bit per round, can be biased by $O(\ell/\log^{(k)}(\ell))$ bad players. For all $k>1$ this strengthens the previous best lower bounds [RSZ, SICOMP 2002], which ruled out protocols resilient to adversaries controlling $O(\ell/\log^{(2k-1)}(\ell))$ players. As a consequence, we establish that any protocol tolerating a linear fraction of corrupt players, while restricting player messages to 1 bit per round, must run for at least $\log^{*}\ell-O(1)$ rounds, improving on the prior best lower bound of $\frac{1}{2}\log^{*}\ell-\log^{*}\log^{*}\ell$. This lower bound also matches the number of rounds, $\log^{*}\ell$, taken by the current best coin flipping protocols from [RZ, JCSS 2001], [F, FOCS 1999] that can handle a linear sized coalition of bad players, given the additional freedom that players can send unlimited bits per round. We also extend our techniques to derive lower bounds for protocols allowing multi-bit messages per round. Our results show that the protocols from [RZ, JCSS 2001], [F, FOCS 1999] that handle a linear number of corrupt players are almost optimal in terms of round complexity and communication per player in a round.
A key technical ingredient in proving our lower bounds is a new result regarding biasing most functions from a family of functions using a common set of bad players and a small specialized set of bad players specific to each function that is biased.
Complementing our lower bound results, we give improved constant-round coin flipping protocols in the setting that each player can send $1$ bit per round. For example, in the case of two rounds, our protocol can handle $O(\ell/(\log\ell)(\log\log\ell)^{2})$ sized coalition of bad players; this is better than the best one-round protocol (also called a resilient function) by [AL, Combinatorica 1993] in this setting, which can only handle $O(\ell/(\log\ell)^{2})$ sized coalition of bad players.
@@ 1 Introduction
Two related and fundamental tasks in fault tolerant distributed computing are for processors to (1) collectively flip a coin, i.e. agree on a common random bit, and to (2) select a ‘leader’ among them, in the presence of adversarial faults. These tasks have been widely studied under various assumptions about the communication channels between the processors, the kind of faults allowed, the power of the adversary, and the kinds of randomness each processor has access to. We study this problem under the assumption that there is only a single common broadcast channel, that adversarial processors are computationally unbounded, and that each processor has access to private randomness. This widely studied model was introduced by Ben-Or and Linial [BL85coin], and is known as the full information model since all processors have access to the same information.
@@
The standard way of modeling how processors coordinate their efforts is through the notion of a protocol. All players (processors henceforth will be called players) agree on a fixed protocol $\pi$ beforehand and execute it, at the end of which they all agree on a common random bit or a leader. We assume that there is an adversary $\mathcal{A}$ that selects a subset of players before the protocol begins and continues controlling them throughout the execution of the protocol. We refer to the coalition of controlled players as ‘bad’ and the remaining players as ‘good’. A protocol consists of one or more rounds. In each round, all players should flip $r$ private random coins and broadcast those $r$ bits to everyone else. The identity of the sender of the bits is always known. This continues for $k$ rounds, after which the protocol determines the common random bit or selects the leader. We assume that each round of the protocol is asynchronous, and we always consider the worst-case scenario in which all good players broadcast their bits at the beginning of the round. Then, based on their outputs, the adversary $\mathcal{A}$ determines the $r$ bits output by each of the bad players. As a result, the outputs of the bad players are coordinated and depend on the outputs of the good players and the outputs of previous rounds. We note that the players are synchronized in between rounds, i.e., a round ends only when all outputs of all players are received. For formal definitions of these protocols, we refer the reader to Section 3.3.1.
@@
Since bad players in this model are computationally unbounded, cryptography-based protocols, which are standard for the Byzantine generals problem and other related models where the adversary is (say) polynomially bounded, are not useful in our setting. Nevertheless, remarkable protocols do exist in this model that can guarantee that a good leader is always chosen with non-trivial probability, or that the outcome of the coin flip is not too biased towards any particular result.
@@ 1.1 Related Work
We first survey the best protocol constructions that are currently known. We mention that given a leader election protocol, one can turn it into a collective coin flipping protocol at the expense of one extra round by making the elected leader flip a coin and output the result (see 3.10 for a proof). Hence, most constructions have focused on leader election protocols. In all these protocols, the goal is to minimize the number of rounds, minimize the number of random bits each player can send per round, and maximize the number of bad players the protocol can handle while electing a good leader with non-trivial probability / flipping a coin where each output in $\{0,1\}$ has at least constant probability of appearing.
@@
Starting with Ajtai and Linial’s non-explicit function [AL93resilient], later made explicit by [Chattopadhyay2019TwoSource, Meka17resilient, IMV23resilient, IV25resilient], one can construct a one-round collective coin flipping protocol, where each player sends $1$ random bit and the output remains close to uniform even in the presence of $O(\ell/(\log\ell)^{2})$ bad players. Leader election protocols that can guarantee a constant probability of electing a good leader in the presence of $O(\ell)$ bad players have been widely studied: when the number of random bits per round is restricted to be $1$, a long line of works [BL85coin, Saks89leader, AN93leader, CL95leader, ORV94leader, Zuckerman97sample, feige99lightestbin, Antonakopoulos06leader] constructed various explicit protocols that run in $O(\log\ell)$ rounds. Moreover, when the number of random bits per round is unrestricted, [RZ01leader, feige99lightestbin] constructed explicit leader election protocols that run in $\log^{*}(\ell)+O(1)$ rounds. [^1] Recall that $\log^{*}(\ell)$ is the minimum number of logs that need to be applied to $\ell$ until it attains value at most $1$.
@@
On the negative side, very few results are known regarding the non-existence of protocols with various parameters. Before stating the known results, we observe that since leader election protocols imply coin flipping protocols, lower bounds for coin flipping protocols imply lower bounds for leader election protocols with one less round (see Corollary 3.11 for formal claim). Hence, all efforts to date have focused on providing lower bounds for coin flipping protocols. First, it is well known that no protocol can handle $\ell/2$ or more bad players [Saks89leader], regardless of the number of rounds and the number of random bits allowed per round. For 1 round coin flipping protocols where players send 1 bit per round, [KKL88influence] showed every protocol can be biased towards some outcome by some set of $O(\ell/\log\ell)$ bad players, almost matching the construction of [AL93resilient]. For multiple rounds, [RSZ02leader] showed that any $k$-round coin flipping protocol where each player sends 1 bit per round can be biased towards some outcome by $O(\ell/\log^{(2k-1)}\ell)$ bad players. [^2] We use the notation $\log^{(i)}\ell$ to denote $i$-times-iterated logarithm, i.e. $\overbrace{\log\cdots\log}^{i\text{~{}times}}\ell$. This also implies that in order to handle $\Theta(\ell)$ bad players when the number of bits per round is $1$, the number of rounds required is $\frac{1}{2}\log^{*}(\ell)-\log^{*}\log^{*}\ell$. For coin flipping protocols where players can send $\left(\log^{(2i-1)}\ell\right)^{1-o(1)}$ random bits in round $i$, [RSZ02leader] showed that there always exists a set of $o(\ell)$ bad players that can bias them towards some outcome.
@@
The work of Filmus et al. [FHHHZ19] gave lower bounds on collective coin flipping protocols under arbitrary product distributions, rather than the uniform distribution, on the Boolean cube. For further details, we refer the reader to the excellent survey of Dodis [dodis2006fault] on protocols and lower bounds in the full information model.
@@ 1.2 Our Results
Our main results are various improved lower bounds for coin flipping protocols. As noted above, since a leader election protocol can be turned into a coin flipping protocol (with one extra round), our lower bounds for coin flipping protocols immediately imply lower bounds for leader election protocols as well.
Our lower bounds improve and subsume all previous multi-round lower bounds (established by [RSZ02leader]). As discussed below, in some sense our lower bounds show that the protocols of [RSZ02leader] and [feige99lightestbin] are optimal.
@@
For any $k$-round coin flipping protocol $\pi$ over $\ell$ players where players send 1 bit per round, there exists a set of bad players $B\subset[\ell]$ and an outcome $o\in\{0,1\}$, with $\left\lvert B\right\rvert\leq O\left(\frac{\ell}{\log^{(k)}(\ell)}\right)$, such that the players in $B$ can bias $\pi$ to output $o$ with probability $0.99$.
This directly improves upon the previous best lower bounds, due to [RSZ02leader], which showed that $O\left(\frac{\ell}{\log^{(2k-1)}(\ell)}\right)$ players can corrupt such protocols, achieving a strictly better bound for all $k>1$.
We also obtain a lower bound on the number of rounds required for protocols that can handle a linear number of bad players:
@@
For any $k$-round coin flipping protocol $\pi$ over $\ell$ players where players send 1 bit per round and where $k\leq\log^{*}\ell-O(1)$, there exists a set of bad players $B\subset[\ell]$ with $\left\lvert B\right\rvert\leq 0.01\ell$ and an outcome $o\in\{0,1\}$ such that the players in $B$ can bias $\pi$ to output $o$ with probability $\geq 0.99$.
This result shows that any coin flipping protocol, where players send 1 bit per round, that can handle linear sized coalitions requires $\log^{*}\ell-O(1)$ rounds. This improves upon the best previous lower bound of $\frac{1}{2}\log^{*}\ell-\log^{*}\log^{*}\ell$ rounds, due to [RSZ02leader] . This lower bound also essentially matches the number of rounds, $\log^{*}\ell$, used by the current best coin flipping protocols from [RZ01leader, feige99lightestbin] that can handle linear sized coalition of bad players, given the additional freedom that players can send unlimited bits per round.
We remark that a key ingredient towards proving 1 is a new lemma regarding biasing most functions from a family of functions using a common set of bad players and a small specialized set of bad players specific to each function being biased. For a formal statement, refer to Theorem 4.1
We also obtain lower bounds against protocols where players can send multiple bits per round:
@@
Let $\pi$ be a $k$-round coin flipping protocol over $\ell$ players where each player can send $(\log^{(i)}\ell)^{0.99}$ many bits in round $i$, and $k\leq\log^{*}\ell-O(1)$. Then, there exists a set of bad players $B\subset[\ell]$ and an outcome $o\in\{0,1\}$, with $\left\lvert\mathbf{B}\right\rvert\leq 0.01\ell$, such that the players in $B$ can bias $\pi$ to output $o$ with probability $\geq 0.99$.
The coin flipping protocols from [RZ01leader, feige99lightestbin] that can handle a linear sized coalition of bad players take $\log^{*}\ell+O(1)$ rounds, and in round $i$, only require that all players send $O(\log^{(i)}\ell)$ many bits. Hence, 2 shows that the protocols from [RZ01leader, feige99lightestbin] are essentially optimal, in the sense that if the number of rounds and the number of bits each player can send in round $i$ are even slightly lowered, then no protocol can handle a linear sized coalition of bad players. This result also directly improves on a bound in this setting obtained by [RSZ02leader]; their result showed that if each player can send at most $(\log^{(2i-1)}\ell)^{0.99}$ bits in round $i$, then any such protocol cannot handle a linear sized coalition of bad players. This resolves an open problem raised in [RSZ02leader], where they exactly asked for a result along the lines of 2.
We also complement the lower bound result described above, by constructing improved constant-round protocols in which each player can send one bit per round. Our goal is to maximize the number of bad players the protocol can tolerate while still ensuring that the output coin flip remains nearly balanced. We provide constructions of such protocols.
@@
For any $k\geq 2$, there exists an explicit $k$-round coin flipping protocol over $\ell$ players where each player sends $1$ bit per round such that when the number of bad players is at most $O\left(\frac{\ell}{(\log\ell)(\log^{(k)}\ell)^{2}}\right)$, the output bit is $0.01$-close to uniform.
@@
When $k=2$, this protocol can handle any $O\left(\frac{\ell}{(\log\ell)(\log\log\ell)^{2}}\right)$-sized coalition of bad players, which is better than the best $1$ round protocol by [AL93resilient] in this setting which can only handle an $O(\ell/(\log\ell)^{2})$-sized coalition of bad players. To compare with our lower bound, when $k=2$ our lower bound from 1 shows any such protocol can be biased by a $O(\ell/(\log\log\ell))$ sized coalition of bad players. It is a very interesting open problem to settle the exact number of bad players that the best two-round coin flipping protocol can handle.

### v2

# Improved Bounds for Coin Flipping, Leader Election, and Random Selection

@@ Abstract
Random selection is a fundamental task in fault-tolerant distributed computing where processors select a random outcome from some domain. Two special cases of this, leader election (where the processors designate a leader amongst themselves) and collective coin flipping (where the processors agree on a common random bit), have been especially widely studied. We study these problems in the full-information model, where processors communicate via a single broadcast channel, have access to private randomness, and face a computationally unbounded adversary that controls some of the processors. Despite decades of study, key gaps remain in our understanding of the trade-offs between round complexity, communication per player in each round, and adversarial resilience. We make progress by proving new lower bounds for coin flipping protocols and both new upper and lower bounds for leader election and random selection protocols.
We first show that any $k$-round coin flipping protocol, where each of $\ell$ players sends $1$ bit per round, can be biased by $O(\ell/\log^{(k)}(\ell))$ bad players. We obtain the same lower bound (with an additional $\log^{(k+1)}(\ell)$ factor in the numerator) for leader election as well. This strengthens the previous best lower bounds [RSZ, SICOMP 2002], which ruled out coin flipping protocols resilient to $O(\ell/\log^{(2k-1)}(\ell))$ bad players and leader election protocols resilient to $O(\ell/\log^{(2k+1)}(\ell))$ bad players. As a consequence, we establish that any protocol tolerating a linear fraction of corrupt players, while restricting player messages to 1 bit per round, must run for at least $\log^{*}\ell-O(1)$ rounds, improving on the prior best lower bound of $\frac{1}{2}\log^{*}\ell-\log^{*}\log^{*}\ell$. We additionally show that the current best protocols that handle a linear number of corrupt players (from [RZ, JCSS 2001], [F, FOCS 1999]) are near optimal in terms of round complexity and communication per player in a round.
We next initiate the study of one-round random selection protocols where each player sends 1 bit in the round. We obtain a one-round protocol resilient to $\ell/(\log(\ell))^{2}$ bad players that outputs $(\log(\ell))^{2}/(\log(\log(\ell)))^{2}$ uniform random bits. As a consequence we obtain a one-round leader election protocol resilient to $\ell/(\log(\ell))^{2}$ bad players, improving on the previous best protocol from [RZ, JCSS 2001] that is resilient to only $\ell/(\log(\ell))^{3}$ bad players and requires players to send many bits. Our resilience parameter matches that of the best one-round coin flipping protocol by Ajtai and Linial, which only outputs one bit. We also obtain an almost matching lower bound: we show that any protocol that outputs $(\log(\ell))^{2}/(\log(\log(\ell)))^{2}$ uniform random bits can be corrupted using $\ell(\log(\log(\ell)))^{2}/(\log(\ell))^{2}$ bad players. In fact, our lower bound provides a more general tradeoff that lets us rule out random selection protocols with far weaker guarantees. To obtain our lower bound, we introduce and study multi-output influence, a natural extension of the notion of influence of boolean functions to the multi-output setting.
@@ 1 Introduction
A fundamental task in fault tolerant distributed computing is random selection, where processors collectively select a random outcome from a domain in the presence of adversarial faults. Two special cases of this, collective coin flipping, where processors agree on a common random bit, and leader election, where the task is to select a “leader” amongst them, are extremely important. These tasks have been widely studied under various assumptions about the communication channels between the processors, the kind of faults allowed, the power of the adversary, and the kinds of randomness each processor has access to. We study this problem under the assumption that there is only a single common broadcast channel, that adversarial processors are computationally unbounded, and that each processor has access to private randomness. This extensively studied model was introduced by Ben-Or and Linial [BL85], and is known as the full information model since all processors have access to the same information.
@@
The standard way of modeling how processors coordinate their efforts is through the notion of a protocol. All players (processors henceforth will be called players) agree on a fixed protocol $\pi$ beforehand and execute it, at the end of which they all agree on an outcome from the output domain. We assume that there is an adversary $\mathcal{A}$ that selects a subset of players before the protocol begins and continues controlling them throughout the execution of the protocol. We refer to the coalition of controlled players as “bad” and the remaining players as “good”. A protocol consists of one or more rounds. In each round, all players should flip $r$ private random coins and broadcast those $r$ bits to everyone else. The identity of the sender of a bit is always known. This continues for $k$ rounds, after which the protocol determines the outcome. We assume that each round of the protocol is asynchronous, and we always consider the worst-case scenario in which all good players broadcast their bits at the beginning of the round. Then, based on their outputs, the adversary $\mathcal{A}$ determines the $r$ bits output by each of the bad players. As a result, the outputs of the bad players are coordinated and depend on the outputs of the good players and the outputs of previous rounds. We note that the players are synchronized in between rounds; i.e., a round ends only when all outputs of all players are received. For formal definitions of these protocols, we refer the reader to Section 3.3.1.
@@
Since bad players in this model are computationally unbounded, cryptography-based protocols, which are standard for the Byzantine generals problem and other related models where the adversary is (say) polynomially bounded, are not useful in our setting. Nevertheless, remarkable protocols do exist in this model that can guarantee that a good leader is always chosen with non-trivial probability, or that the outcome of the collective coin flip has small bias.
When constructing random selection protocols, the goal is to minimize the number of rounds, minimize the number of bits each player sends per round, and maximize the number of bad players the protocol can handle, while attaining the objective of making the selected outcome as “random” as possible. The ideal objective in random selection is that (even in the presence of bad players), the output distribution is statistically very close (or even just non-trivially close) to the uniform distribution over the output domain. When the output domain is $\{0,1\}$, this exactly corresponds to the objective in collective coin flipping. A slightly weaker objective in random selection is that the output distribution should not be too concentrated on a too-small set of outcomes from the domain. This corresponds to the objective in leader election where the task is to elect a good player as a leader with non-trivial probability.  [^1] Concretely, if $b$ out of $\ell$] players are bad then the task is to ensure that the output distribution is not concentrated on any $S\subset[\ell]$ with $\left\lvert S\right\rvert\leq b$.
@@ 1.1 Related Work
We first survey the best protocol constructions that were known prior to our work.
@@
We first mention that given a leader election protocol, one can turn it into a collective coin flipping protocol at the expense of one extra round by making the elected leader flip a coin and output the result (see 3.11 for a proof). Starting with Ajtai and Linial’s non-explicit function [AL93], later made explicit by [CZ19, Mek17, IMV23, IV24], one can construct a one-round collective coin flipping protocol where each player sends $1$ random bit that is resilient to $O(\ell/(\log(\ell))^{2})$ bad players. Leader election protocols that can guarantee a constant probability of electing a good leader in the presence of $O(\ell)$ bad players have also been widely studied: when the number of random bits per round is restricted to be $1$, a long line of works [BL85, Sak89, AN93, CL95, ORV94, Zuc97, BN00, Fei99, Ant06] constructed various explicit protocols, the best of which runs in $O(\log(\ell))$ rounds. Moreover, when the number of random bits per round is unrestricted, [RZ01, Fei99] constructed explicit leader election protocols that run in $\log^{*}(\ell)+O(1)$ rounds. [^2] Recall that $\log^{*}(\ell)$ is the minimum number of logs that need to be applied to $\ell$ until it attains value at most $1$.
The study of general random selection protocols with the weaker objective of not being too concentrated on a small set of outputs was initiated by [GGL98], and later works constructed protocols with stronger guarantees and reduced complexity [SV08, GVZ06]. All these protocols last for many rounds, each player can send many bits per round, and the number of bad players considered is linear in $\ell$ and sometimes even more than $\ell/2$. The work of [KZ07] constructed 1 round random selection protocol with all players sending $1$ bit that is resilient to $b$ bad players and outputs $\ell/(b^{1.58})$ random bits.
@@
On the lower bound side, very few results are known and they have focused mainly on collective coin flipping since lower bounds here translate into lower bounds for leader election protocols with one fewer round (see Corollary 3.12 for a formal claim). First, it is well known that no coin flipping or leader election protocol can handle $\ell/2$ or more bad players [Sak89], regardless of the number of rounds and the number of random bits allowed per round. For one-round coin flipping protocols where players send 1 bit per round, [KKL88] showed that every protocol can be biased towards some outcome by some set of $O(\ell/\log(\ell))$ bad players. For multiple rounds, [RSZ02] showed that any $k$-round coin flipping protocol where each player sends 1 bit per round can be biased towards some outcome by $O(\ell/\log^{(2k-1)}(\ell))$ bad players. [^3] We use the notation $\log^{(i)}\ell$ to denote $i$-times-iterated logarithm, i.e. $\overbrace{\log\cdots\log}^{i\text{~times}}\ell$. This also implies that in order to handle $\Theta(\ell)$ bad players when the number of bits per round is $1$, the number of rounds required is $\frac{1}{2}\log^{*}(\ell)-\log^{*}(\log^{*}(\ell))$. For random selection tasks, [GGL98] showed that for any $\mu>0$, $b$ bad players can always ensure that the output lies in some set of density $\mu$ with probability at least $\mu^{1-(b/\ell)}$. Kamp and Zuckerman showed that for any one round random selection protocol outputting $m$ bits and where all players send $1$ bit, $O\left(\frac{\ell}{m}\right)$ bad players can corrupt the protocol so that the output distribution is $0.1$-far from $\mathbf{U}_{m}$ [KZ07].
@@
The work of [FHHHZ19] gave lower bounds on collective coin flipping protocols under arbitrary product distributions, rather than the uniform distribution, on the Boolean cube. For further details, we refer the reader to the excellent survey of Dodis [Dod06] on protocols and lower bounds in the full information model.
We mention that one-round coin flipping protocols, random selection protocols that output many uniform random bits, and random selection protocols where the output is not too concentrated on a small set, naturally arise in extractor theory where these protocols are referred to as extractors / condensers for Non-Oblivious Symbol Fixing (NOSF) and Non-Oblivious Bit Fixing (NOBF) sources. These protocols have been a pivotal ingredient in a surge of works over the past decade that have resulted in constructions of near optimal two source extractors, two source condensers, affine extractors, sumset extractors, and more [CZ19, Li16, BDT17, BCDT19, CGL21, CL22, Li23].
@@ 1.2 Our Results
We broadly obtain two kinds of results: (1) Improved lower bounds for coin flipping protocols and (2) Improved upper and lower bounds for one-round random selection protocols. For ease of exposition we will specialize and also describe our results in strand (2) for leader election protocols.
1.2.1 Improved lower bounds for collective coin flipping
Our lower bounds for collective coin flipping improve and subsume all previous $k$ round lower bounds (established by [RSZ02]) for $k>1$.
Our first non-trivial improvement is for two-round protocols where players send $1$ bit per round. We show that $\frac{\ell}{\log(\log(\ell))}$ players suffice to corrupt any such protocol. Since the previous best bound was $\frac{\ell}{\log(\log(\log(\ell)))}$, due to [RSZ02], we thus exponentially improve the bound on the fraction of players that can corrupt such a protocol.  [^4] We recall that for one-round protocols, there is a multiplicative gap of $\log(\ell)$ between the upper bound of $\frac{\ell}{(\log(\ell))^{2}}$ and the lower bound of $\frac{\ell}{\log(\ell)}$. Resolving this is a major question in Boolean Fourier analysis. Generally, we obtain:
@@ 1.2.1 Improved lower bounds for collective coin flipping
Let $\pi$ be a $k$-round coin flipping protocol over $\ell$ players where players send 1 bit per round. Then, there exists an outcome $o\in\{0,1\}$ such that $b=O\left(\frac{\ell}{\log^{(k)}(\ell)}\right)$ bad players can bias $\pi$ to output $o$ with probability $\geq 0.99$.
For all $k>1$, this directly improves upon the previous best lower bounds, due to [RSZ02], which showed that $O\left(\frac{\ell}{\log^{(2k-1)}(\ell)}\right)$ players can corrupt such protocols.
Arguably, the most important open problem in collective coin flipping and leader election is to figure out the optimal number of rounds for a protocol handling linear fraction of bad players. The above result lets us make progress on this question by giving us the following lower bound:
@@
Let $\pi$ be a $k$-round coin flipping protocol over $\ell$ players where players send 1 bit per round and where $k\leq\log^{*}(\ell)-O(1)$. Then, there exists an outcome $o\in\{0,1\}$ such that $b=0.01\ell$ players can bias $\pi$ to output $o$ with probability $\geq 0.99$.  [^5] Since $k$-round leader election protocols imply ($k+1$)-round coin flipping protocols, this lower bound also applies to leader election protocols.
This shows that any coin flipping protocol that can handle linear sized coalitions, where players send $1$ bit per round, requires at least $\log^{*}(\ell)-O(1)$ rounds. This improves upon the best previous lower bound of $\frac{1}{2}\log^{*}(\ell)-\log^{*}(\log^{*}(\ell))$ rounds due to [RSZ02].
Our lower bound also essentially matches the number of rounds, $\log^{*}(\ell)$, used by the current best coin flipping protocols from [RZ01, Fei99] that can handle linear sized coalition of bad players, given the additional freedom that players can send unlimited bits per round. This raises the natural question of whether we can further meaningfully show that the protocols of [RZ01, Fei99] are optimal. We indeed show this by obtaining the following lower bound on protocols where players can send many bits per round:
@@
Let $\pi$ be a $k$-round coin flipping protocol over $\ell$ players where each player can send $(\log^{(i)}(\ell))^{0.99}$ many bits in round $i$, and $k\leq\log^{*}(\ell)-O(1)$. Then, there exists an outcome $o\in\{0,1\}$ such that $b=0.01\ell$ players can bias $\pi$ to output $o$ with probability $\geq 0.99$.  [^6] Similarly, this lower bound also applies to leader election protocols.
The coin flipping protocols from [RZ01, Fei99] that can handle a linear sized coalition of bad players and take $\log^{*}(\ell)+O(1)$ rounds have the following property that in round $i$, they only require that all players send $O(\log^{(i)}(\ell))$ many bits. Hence, 2 shows that the protocols from [RZ01, Fei99] are essentially optimal, in the sense that if the number of rounds and the number of bits each player can send in round $i$ are even slightly lowered, then no protocol can handle a linear sized coalition of bad players.
A weaker result of this flavor was obtained by [RSZ02]; their result showed that if each player can send at most $(\log^{(2i-1)}(\ell))^{0.99}$ bits in round $i$, then any such protocol cannot handle a linear sized coalition of bad players. 2 in fact resolves an open problem raised in [RSZ02] where they had exactly asked for the result that we obtain.
We remark that a key ingredient towards proving 1 is a new lemma regarding biasing most functions from a family of functions using a common set of bad players and a small specialized set of bad players specific to each function being biased. For a formal statement, refer to Theorem 4.1.
We complement the lower bound result described above by constructing improved constant-round protocols in which each player can send one bit per round. We obtain:
@@
For any $k\geq 2$, there exists an explicit $k$-round coin flipping protocol over $\ell$ players where each player sends $1$ bit per round such that when the number of bad players is at most $O\left(\frac{\ell}{\log(\ell)\cdot(\log^{(k)}(\ell))^{2}}\right)$, the output bit is $0.01$-close to the uniform distribution.
@@
When $k=2$, this protocol can handle any $O\left(\frac{\ell}{\log(\ell)\cdot(\log(\log(\ell)))^{2}}\right)$-sized coalition of bad players, which is better than the best one-round protocol by [AL93] in this setting that is resilient to only a $O(\ell/(\log(\ell))^{2})$-sized coalition of bad players.
1.2.2 Improved bounds for leader election
We obtain improved upper and lower bounds for leader election protocols. Beginning with our lower bounds, our results subsume all previous lower bounds for multi-round leader election protocols. Our first non-trivial result is for one-round protocols where the players send $1$ bit per round. We show that $\frac{\ell\log(\log(\ell))}{\log(\ell)}$ players suffice to corrupt any such protocol. This is a double exponential improvement upon the previous best lower bound of $\frac{\ell}{\log(\log(\log(\ell)))}$ due to [RSZ02]. [^7] Recall that given a leader election protocol, one can turn it into a collective coin flipping protocol at the expense of one extra round. So $k$-round collective coin flipping lower bounds translate to (k1k-1)-round leader election lower bounds. This was the exact strategy followed by [RSZ02]. We obtain the following general lower bound for $k$-round protocols:
Theorem 4 (Informal version of Corollary 5.3).
Let $\pi$ be a $k$-round leader election protocol over $\ell$ players where players send 1 bit per round. Then, there exist $b=O\left(\frac{\ell\log^{(k+1)}(\ell)}{\log^{(k)}(\ell)}\right)$ bad players who can bias $\pi$ to elect a bad player as a leader with probability $0.99$.
This directly improves upon the previous best lower bounds, due to [RSZ02], which showed that $O\left(\frac{\ell}{\log^{(2k+1)}(\ell)}\right)$ players can corrupt such protocols, achieving a strictly better bound for all $k$. For other lower bounds, we recall that our lower bounds from 1 and 2 also yield leader election lower bounds.
Turning to positive results, we construct the following one-round leader election protocol.
Theorem 5 (Informal version of Corollary 7.3 ).
There exists an explicit one-round leader election protocol $\pi$ over $\ell$ players where players send 1 bit per round such that the following holds: For any set of $b$ bad players with $b\leq O\left(\frac{\ell}{(\log(\ell))^{2}}\right)$, the protocol elects a good leader with probability $0.99$.
The $O\left(\frac{\ell}{(\log(\ell))^{2}}\right)$ number of bad players that this protocol can handle is exactly the number of bad players that the current best one-round coin flipping protocol by Ajtai and Linial can handle. Previously, the best one-round leader election protocol, constructed by [RZ01], required each player to send $O(\log(\ell))$ bits, and even then the protocol could only handle $O\left(\frac{\ell}{(\log(\ell))^{3}}\right)$ bad players.
1.2.3 Bounds for one-round random selection
Here we construct protocols and provide lower bounds for random selection in the natural setting of one-round protocols where each player sends $1$ bit per round.
We construct the following protocols for this problem. Our protocols will meet the gold standard for random selection—the output distribution will be statistically close to the uniform distribution.
Theorem 6 (Informal version of Theorem 7.1).
Let $b,\ell$ be such that $b\leq O\left(\frac{\ell}{(\log(\ell))^{2}}\right)$. Then, there exists an explicit one-round random selection protocol $\pi:\{0,1\}^{\ell}\to\{0,1\}^{m}$ where $m=O\left(\frac{\ell}{b\cdot\left(\log(\ell/b)\right)^{2}}\right)$ such that for any coalition of $b$ bad players corrupting $\pi$, the output distribution on $m$ bits has statistical distance at most $\varepsilon$ from $\mathbf{U}_{m}$.
This improves upon the previous best construction of Kamp and Zuckerman in this setting that outputted $m=O\left(\frac{\ell}{b^{\log(3)}}\right)$ uniform random bits [KZ07].
To get a sense of the parameters, we see that when $b=O\left(\frac{\ell}{(\log(\ell))^{2}}\right)$, we obtain $\left(\frac{\log(\ell)}{\log(\log(\ell))}\right)^{2}$ truly random bits. This also directly gives us a leader election protocol as in 5. This “resilience” parameter of $b=O\left(\frac{\ell}{(\log(\ell))^{2}}\right)$ is exactly the resilience parameter obtained by the current best coin flipping protocol by Ajtai and Linial [AL93]. Hence, we improve upon their result by outputting many more bits while maintaining the resilience parameter.
We obtain the following lower bounds on random selection (even with weaker objectives):
Theorem 7 (Informal version of Theorem 8.1).
For any one-round random selection protocol $\pi:\{0,1\}^{\ell}\to\{0,1\}^{m}$ and any $\Delta>0$, there exists a set of bad players $B\subseteq[\ell]$ with $\left\lvert B\right\rvert=O\left(\frac{\ell}{m}\cdot(\Delta\log(\Delta))^{2}\right)$ and a small set $S\subseteq\{0,1\}^{m}$ with $\left\lvert S\right\rvert\leq 2^{m-\Delta}$ such that the players in $B$ can corrupt $\pi$ to make the protocol output an element of $S$ with probability at least $0.99$.
When $\Delta=O(1)$, the above lower bound rules out protocols outputting near random bits.
Corollary 2.
For any $\varepsilon>0$ and any one-round random selection protocol $\pi:\{0,1\}^{\ell}\to\{0,1\}^{m}$, there exists a set of bad players $B\subseteq[\ell]$ with $\left\lvert B\right\rvert=O\left(\frac{\ell}{m}\right)$ such that the players in $B$ can corrupt $\pi$ so that the output distribution of the protocol has statistical distance at least $1-\varepsilon$ from the uniform distribution.
Previous lower bounds in this setting were obtained by [KZ07] that showed $O\left(\frac{\ell}{m}\right)$ bad players can corrupt the protocol so that the output distribution has statistical distance at least $0.1$ from the uniform distribution. From their techniques, it is not clear how to obtain stronger result that rules out larger statistical distance or stronger random selection impossibility results. Ruling out such protocols is important since in the collective coin flipping literature, even protocols attaining statistical distance at most $0.9$ are considered good.
Once again to get a sense of the parameters, we see that when $m=\left(\frac{\log(\ell)}{\log(\log(\ell))}\right)^{2}$, then $b=O\left(\frac{\ell}{(\log(\ell))^{2}}\cdot(\log(\log(\ell)))^{2}\right)$ bad players can corrupt a protocol to have large statistical distance from the uniform distribution. Hence, these results shows that our one-round random selection protocols are tight for this value of $m$ up to $\operatorname{poly}(\log(\log))$ factors. This is quite surprising since in one-round leader election as well as in collective coin flipping, the multiplicative gap in the number of bad players between the lower bounds and the best known construction is at least $O(\log(\ell))$.
One key definition that we introduce to help prove our lower bounds is that of the multi-output influence of a Boolean function. This naturally generalizes the standard notion of influence of a Boolean function to larger output domains. See Definition 2.18 for the formal definition, our bounds on it, and how we use it.